OpenStudy (anonymous):

Suppose your police department is planning the purchase of 70 new laptops for its patrol cars. Research and find a particular selling price for any style of laptop. Suppose that the laptops are marked down 3% off the total price. Calculate how much the department will save with the 3% savings. What might be reasons that the laptop provider would give the police department a markdown on the purchase of those laptops? Explain clearly.

OpenStudy (anonymous):

If you do answer my question please explain i need to have each step

OpenStudy (anonymous):

it would save .3(price)(70 laptops) because 1) they buy in bulk 2) it's a public department, just as a donation 3) because they have a contract with the company

OpenStudy (anonymous):

So would you multiply use the price 480 and re explain it please

OpenStudy (amistre64):

the math is simple enough, the other parts tho are just something for you to consider

OpenStudy (anonymous):

Yea but when i try to deduct the 3% it says 14.4 i need and exact answer so could you help

OpenStudy (amistre64):

look online for the price of a laptop; and use that for the "price" part

OpenStudy (amistre64):

whats your laptop price?

OpenStudy (anonymous):

480

OpenStudy (amistre64):

ok, 480 * 30% = savings 480/100 = 1%; 3 of those 1% is 3% 3*480/100 = savings

OpenStudy (amistre64):

14.4 is an exact number, but for money we like to see it in 2 deci spots; 14.40

OpenStudy (amistre64):

so, we save $14.40 on each laptop and we have 70 of them to buy so out total savings is? 70 * 14.40

OpenStudy (anonymous):

i dont understand the first part could you explain how you got that

OpenStudy (amistre64):

i calculate that to; $1,008.00

OpenStudy (amistre64):

how I got what exactly?

OpenStudy (amistre64):

precent means: per 100. /100

OpenStudy (amistre64):

percent even, gotta read thru the typos

OpenStudy (amistre64):

1% = total/100

OpenStudy (amistre64):

once we know 1% we can add up as many as we need to get to any % we want